OpenStudy (anonymous):

Chance can wash the car in 2 hours, while Sandy can wash the same car in 4 hours. How many hours would it take for them to wash the car together?

OpenStudy (imstuck):

This is a work problem. If Chance can wash a car in 2 hours, he can get 1/2 the job done in 1 hour. If Sandy can wash the car in 4 hours, she can get 1/4 of the job done in 1 hour. So your equation looks like this:

OpenStudy (imstuck):

\[\frac{ 1 }{ 2 }+\frac{ 1 }{ 4 }=\frac{ 1 }{ x }\]That 1/x is the amount of time it will take them to get the whole job done.

OpenStudy (imstuck):

Find the least common multiple to rid yourself of the denominators and solve for x. 4x is the LCM. Like this:

OpenStudy (imstuck):